OpenStudy (anonymous):

A pharmacist mixed some 10%-saline solution with some 15%-saline solution to obtain 100 mL of a 12%-saline solution. How much of the 10%-saline solution did the pharmacist use in the mixture? A 60 mL B 45 mL C 40 mL D 25 mL

OpenStudy (anonymous):

put \(x\) = ml of 15% solutions and since the total is 100 the amount of 10% solution is \(100-x\) therere the total amount of saline you have is \(.15x+.10(100-x)\) and you want this to be equal to 12% of 100 = 12 so solve \[.15x+.10(100-x)=12\] for \(x\)

OpenStudy (anonymous):

maybe easier to solve \[15x+10(100-x)=1200\] \[15x+1000-10x=1200\] \[5x=200\] etc

OpenStudy (anonymous):

the answer should be 60 ml...so how do I get it? in your solution the answer is 40 ml...